FLOSS01: Logbook Entries

FLOSS01: Status Messages: 30 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
174 Thu 14-Mar-2002Daily status, 3/14/02semmer
167 Tue 12-Mar-2002Daily status, 3/12/02semmer
163 Mon 04-Mar-2002Sta 3 Outages 3/2, 3/3stn3militzer
160 Wed 27-Feb-2002Guy tension/chem sheltertowerdelany
151 Thu 21-Feb-2002Guy wire tension adjustmenttowerdelany
133 Tue 12-Feb-2002all stnsdelany
132 Sat 09-Feb-2002Weekly Status, 2-9 Feb 2002militzer
130 Fri 25-Jan-2002Daily status, 1/25/02semmer
129 Thu 24-Jan-2002Daily status, 1/24/02semmer
128 Wed 23-Jan-2002Daily status, 1/23/02semmer
127 Wed 23-Jan-2002Daily status, 1/22/02semmer
126 Tue 22-Jan-2002Daily status, 1/18/02horst
125 Tue 22-Jan-2002Daily status, 1/17/02horst
121 Fri 18-Jan-2002Installed web camtowermaclean
114 Wed 16-Jan-2002Daily status, 1/16/02horst
113 Tue 15-Jan-2002Daily status, 1/15/02horst
112 Mon 14-Jan-2002Daily Status, 1/14/02horst
79 Sun 09-Dec-2001visitorsall stnsoncley
76 Sat 08-Dec-2001shallow inversion todayall stnsoncley
70 Sat 08-Dec-2001serial numbers and heightsall stnsoncley
51 Wed 05-Dec-2001site visittoweroncley
47 Tue 04-Dec-2001second mouse trappedbaseoncley
37 Mon 03-Dec-2001covars restartedbaseoncley
31 Mon 03-Dec-2001miscstn2oncley
29 Mon 03-Dec-2001daily statusall stnsoncley
25 Sun 02-Dec-2001Dec 2 morning statusall stnsmaclean
22 Sat 01-Dec-2001Operational!all stnsmaclean
14 Thu 29-Nov-2001statusall stnsoncley
10 Thu 29-Nov-2001Tower heightstoweroncley
9 Wed 28-Nov-2001statustoweroncley


174: Status, Site , Thu 14-Mar-2002 12:39:36 MST, Daily status, 3/14/02
FLOSS daily status, 03/14/02

This status is based on QC web plots.

Looked at past 48 hours of data
Webcam shows snow was melting till last night. Now
there is a fresh cover of snow.

Meteorology
-----------
Prop Spd and Dir:	ok, prop at 20m replaced yesterday
T/RH (ISFF):		ok, 5 meter T questionable from 3/13/02 17:00 to 5:00
Pressure:		ok
Rainr:			ok, activity throughout last night
Depth:			ok, corresponds with rain activity
Rsw:			ok
Rlw:			ok
Rnet:			ok

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok













167: Status, Site , Tue 12-Mar-2002 08:28:27 MST, Daily status, 3/12/02
FLOSS daily status, 03/12/02

This status is based on QC web plots.

Looked at past 48 hours of data
Webcam shows snow is melting

Meteorology
-----------
Prop Spd and Dir:	prop at 20m out
T/RH (ISFF):		ok
Pressure:		ok
Rainr:			ok
Depth:			ok
Rsw:			ok
Rlw:			ok
Rnet:			ok

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok












163: Status, Site stn3, Mon 04-Mar-2002 10:10:12 MST, Sta 3 Outages 3/2, 3/3
Station 3 Power Outages

It appears Sage site is getting cold feet.....

---------------------------------

Sta3 - Sage Outage:
3/2/02 ~11:30 - 15:00z
Low voltage disconnect did it's thing.
Station recovered by itself with adequate solar charging,
no wind charging at the moment.

4-Component FANS Disconnected:  ~16:30z
to reduce the load.  They're still jumping up and down in
power consumption when the temp. drops too low.

---------------------------------

Sta3 - Sage Outage:
3/3/02 ~9:15 - 14:45z
Low voltage disconnect did it's thing again.
Station recovered by itself with adequate solar charging.
At about local noon, the winds picked up and the wind
generator was providing charging through the night.



160: Status, Site tower, Wed 27-Feb-2002 16:16:33 MST, Guy tension/chem shelter
Checked guy tensions: see record # 109
Chem shelter moved at 14:30 to position near the entry gate, to remove it
from the upwind fetch of the precip guage.
151: Status, Site tower, Thu 21-Feb-2002 14:54:57 MST, Guy wire tension adjustment
Between, 13:30 and 14:30 I adjusted the tower guy tensions.
First I monitored the 12 individual guys. Then I increased the tension on the 
lowest value guy by one notch. This increased the value for that guy by a 
several hundred pounds. It also modulated the tensions on the other guys.
I monitored all the guys again and then increaded the tension on the new
lowest value guy. I reiterated three times until all tensions were above 400lbs.

See logbook entry 109 for the record.

133: Status, Site all stns, Tue 12-Feb-2002 14:44:22 MST,
FLOSS status, 2002 Feb 11

NOTE: I have just started the data minding task


Webcam:


        Meteorology
        -----------
        Prop Spd/ Dir:  ok     
        T/RH (ISFF):    ok
        Pressure:       ok
        Rainr:



        Depth:          ok value varies
        Rsw:            ok the albedo at the tower is ~60%, at sage only 30%
        Rlw:            ok
        Rnet:           ok different albedos

        Soil
        ----

        Tsoil:          ok, tower still noticably warmer
        Gsoil:          ok, sage still has significantly less flux
        Qsoil:          ok, sage still driest


                        Sonics
        ------
        flags/diag              ok, flag and diag hits on tower 2m.  This appears
                        correlated with temps below -20degC.
        spd, dir, w:            ok the difference plot now fixed        
        variances:              ok the differnce plot now fixed
        u*, sigma_w/u*, w'tc':  ok


        Scalars
        -------

        H2O             ok, The open path Licor is strange beast, probably a calib.
                        issue.  It apparently has inverse trend and
                        gain factor of about 12x compared with kryptons.
                        Sage/Bare still have irregularly high signals
                        at times and these correlate with very low temps
                        going below ~-22degC. 
        co2:            ok still showing calibration peaks, ie still got cal gas
        w'h2o':         ok
        w'co2':         ok


        Status
        ------
        PAM power
        X/Ylevel
        Ice    

        <><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><>



132: Status, Site , Sat 09-Feb-2002 10:43:20 MST, Weekly Status, 2-9 Feb 2002
FLOSS status, 2/9/02

NOTE: I have been looking at data for the last few days.
      This is a combined simplistic status report for the
      last week as opposed to a daily report.

Webcam still has low-resolution, grainy images but appears
	to show good snow cover although it's impossible to
	tell if it's changed much over the period.

Meteorology
-----------
Prop Spd/ Dir:	ok
T/RH (ISFF):	ok
Pressure:	ok, offset in tower analog transducer, calib.?
Rainr:		ok?, many tips with diurnal cycle apparently occuring
			below -10degC.  The sum for 7 days would have been
			1.58" if we can believe that.  There was some snow 'in
			the high country' reported this week.
Depth:		ok, still has diurnal cycle, but has gone up about 10mm
Rsw:		ok, sage.out not as high as tower.out
Rlw:		ok
Rnet:		ok, tower and bare very low values, maybe snow on top
			(but hasn't melted...?), or dirty, etc.  Sage ok.

Soil
----
Tsoil:		ok, tower still noticably warmer (1-deg), cal/instl.?
Gsoil:		ok, sage still has significantly less flux
Qsoil:		ok, sage still driest

Sonics
------
flags/diag		ok, flag and diag hits on tower 2m.  This appears
			correlated with temps below -20degC.
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
h2o:		ok, The open path Licor is strange beast, probably a calib.
			issue.  It apparently has inverse trend and
			gain factor of about 12x compared with kryptons.
		Sage/Bare with serializer still have irregularly high signals
			at times and these correlate with very low temps
			going below ~-22degC.  Could be a calibration issue
			on the serializer a/d.   Only raw data counts reported
			so we're not taking advantage of integrated sensor
			approach.
co2:		ok
w'h2o':		ok
w'co2':		ok


Status
------
PAM power	ok, both stations batt staying up nicely
X/Ylevel	ok, diurnal cycle on tower sensor, X as high as -1deg
Ice		ok?  Got a hit on 5th at about 20:00-20:30 so probably real.














130: Status, Site , Fri 25-Jan-2002 07:18:41 MST, Daily status, 1/25/02
FLOSS daily status, 1/25/02

Looked at past 48 hours of data
Webcam shows snowfall over surface

Meteorology
-----------
Prop Spd and Dir:	ok, winds around 200 degrees
T/RH (ISFF):		ok
Pressure:		ok
Rainr:			ok, slow down in tips
Depth:			ok, has gone up, maybe snow melt
Rsw:			ok, sage.out not as high as tower.out
Rlw:			ok
Rnet:			ok, tower and bare low values, maybe snow on top?

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok











129: Status, Site , Thu 24-Jan-2002 07:18:48 MST, Daily status, 1/24/02
FLOSS daily status, 1/24/02

Looked at past 48 hours of data
Webcam shows snowfall over surface

Meteorology
-----------
Prop Spd and Dir:	ok, winds around 200 degrees
T/RH (ISFF):		ok
Pressure:		ok
Rainr:			ok, a lot of tips during day 23
Depth:			ok, continued to move down
Rsw:			ok, sage.out not as high as tower.out
Rlw:			ok
Rnet:			ok

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok










128: Status, Site , Wed 23-Jan-2002 08:22:35 MST, Daily status, 1/23/02
FLOSS daily status, 1/23/02

Looked at past 48 hours of data
Webcam shows showfall started late yesterday

Meteorology
-----------
Prop Spd and Dir:	ok, winds 150-200 degrees
T/RH (ISFF):		ok
Pressure:		ok
Rainr:			ok, a lot of tips from last night to this morning
Depth:			ok, change started at midnight. Corresponds with
			    rainr info. Webcam confirms snowfall
Rsw:			ok, sage.out not as high as tower.out
Rlw:			ok
Rnet:			ok

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok









127: Status, Site , Wed 23-Jan-2002 08:19:15 MST, Daily status, 1/22/02
FLOSS daily status, 1/22/02

Looked at past 72 hours of data


Meteorology
-----------
Prop Spd and Dir:	ok, winds 150-250 degrees
T/RH (ISFF):		ok
Pressure:		ok
Rainr:			ok, some tips
Depth:			ok, spiking
Rsw:			ok
Rlw:			ok
Rnet:			ok

Soil
----
Tsoil:			ok
Gsoil:			ok
Qsoil:			ok

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			ok
co2:			ok
h2o:			ok






126: Status, Site , Tue 22-Jan-2002 10:25:01 MST, Daily status, 1/18/02
FLOSS daily status, 1/18/02

Looked at past 72 hours of data
Comments:  One has to wonder about the continuous sporadic occurrence of
           single tips on the rain gauge.
           Entered into logbook on 1/22/02, since logbook not available 
           on 1/18/02.

Meteorology
-----------
Prop Spd and Dir:	ok, winds mostly 180-270 degrees
T/RH (ISFF):		ok; did not examine thermocouple data
Pressure:		tower 0.5-3 mb higher than bare/sage;
			  sage 0.1 mb higher than bare
Rainr:			ok(?); netcdf file has units of tips/60
Depth:			"snow depth" ~ 2050 mm; difficult to interpret
Rsw:			ok
Rlw:			Rlw.in.tower > Rlw.in.sage by ~5 W/m^2
Rnet:			ok; Rnet < Rsum by 5-10%

Soil
----
Tsoil:			sage has a little larger diurnal cycle than
			  tower or bare
Gsoil:			sage has smallest diurnal cycle, bare greatest
Qsoil:			sage driest, bare wettest; tower very constant

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			not much correlation evident between 
			  w'co2'(LI7500) and w'co2mr'(LI6251);
  			  LI7500 has diurnal cycle; LI6251 much noisier
co2:			good correlation between co2(LI7500) and 
                          co2mr(LI6251); both are anti-correlated with
			  temperature


125: Status, Site , Tue 22-Jan-2002 10:22:56 MST, Daily status, 1/17/02
FLOSS daily status, 1/17/02

Looked at past 72 hours of data
Comments:  Fixed "overrange" problem for KH2O @ sage; we were using the
wrong calibration; Gordon reran covars.
Entered into logbook on 1/22/02, since logbook was not available on
1/17/02.

Meteorology
-----------
Prop Spd and Dir:	ok, winds mostly 180-270 degrees
T/RH (ISFF):		ok; did not examine thermocouple data
Pressure:		tower 0.5-3 mb higher than bare/sage;
			  sage 0.1 mb higher than bare
Rainr:			ok; netcdf file has units of tips/60
Depth:			"snow depth" ~ 2050 mm; difficult to interpret
Rsw:			ok
Rlw:			Rlw.in.tower > Rlw.in.sage by ~5 W/m^2
Rnet:			ok; Rnet < Rsum by 5-10%

Soil
----
Tsoil:			sage has a little larger diurnal cycle than
			  tower or bare
Gsoil:			sage has smallest diurnal cycle, bare greatest
Qsoil:			sage driest, bare wettest; tower very constant

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			ok
w'co2':			not much correlation evident between 
			  w'co2'(LI7500) and w'co2mr'(LI6251);
  			  LI7500 has diurnal cycle; LI6251 much noisier
co2:			good correlation between co2(LI7500) and 
                          co2mr(LI6251); both are anti-correlated with
			  temperature


121: Status, Site tower, Fri 18-Jan-2002 08:40:35 MST, Installed web cam
We installed the axis web camera on the tower just above the 20m prop vane.


Right now this is the configuration for storing photos:

Between 05:00 and 20:00 MST it will store 704x640 images
on /backup/aster/projects/FLOSS01/webcam, every 15 minutes.


There is a problem however: it is truncating the images, only
56000 bytes are being ftp'd, and so we're only getting about
2/3rds of each image.  I've sent an email inquiry to axis.com.







114: Status, Site , Wed 16-Jan-2002 12:19:19 MST, Daily status, 1/16/02
FLOSS daily status, 1/16/02

Looked at past 72 hours of data
Comments:  snow depth difficult to interpret.  KH2O @ sage overrange.

Meteorology
-----------
Prop Spd and Dir:	ok, winds mostly 180-270 degrees
T/RH (ISFF):		ok; did not examine thermocouple data
Pressure:		tower 0.5-3 mb higher than bare/sage;
			  sage 0.1 mb higher than bare
Rainr:			ok; netcdf file has units of tips/60
Depth:			"snow depth" ~ 2050 mm; difficult to interpret
Rsw:			ok
Rlw:			Rlw.in.tower > Rlw.in.sage by ~5 W/m^2
Rnet:			ok; Rnet < Rsum by 5-10%

Soil
----
Tsoil:			sage has a little larger diurnal cycle than
			  tower or bare
Gsoil:			sage has smallest diurnal cycle, bare greatest
Qsoil:			sage driest, bare wettest

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			KH2O at sage appears to be over range
 			  LI7500 ok
w'co2':			not much correlation evident between 
			  w'co2'(LI7500) and w'co2mr'(LI6251);
  			  LI7500 has diurnal cycle; LI6251 much noisier
co2:			good correlation between co2(LI7500) and 
                          co2mr(LI6251); both are anti-correlated with
			  temperature


113: Status, Site , Tue 15-Jan-2002 10:24:42 MST, Daily status, 1/15/02
FLOSS daily status, 1/15/02

Looked at past 72 hours of data
Comments:  snow depth difficult to interpret.  KH2O @ sage died.

Meteorology
-----------
Prop Spd and Dir:	ok, winds mostly 180-270 degrees
T/RH (ISFF):		ok; did not examine thermocouple data
Pressure:		tower 0.5-3 mb higher than bare/sage;
			  sage 0.1 mb higher than bare
Rainr:			ok
Depth:			"snow depth" ~ 2050 mm; difficult to interpret
Rsw:			tower & sage only; albedo decreasing at tower
Rlw:			tower & sage only; ok
Rnet:			ok; Rnet < Rsum by 5-10%

Soil
----
Tsoil:			sage has a little larger diurnal cycle than
			  tower or bare
Gsoil:			sage has smallest diurnal cycle, bare greatest
Qsoil:			sage driest, bare wettest

Sonics
------
spd, dir, w:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			KH2O at sage appears to have died around 1000,
			  Jan 14, over 2 hrs since last precip. 
 			  LI7500 ok
w'co2':			not much correlation evident between w'co2' 
			  from LI7500 and w'co2mr' from LI6251;
  			  LI7500 has diurnal cycle; LI6251 much noisier
co2:			good correlation between co2(LI7500) and 
                          co2mr(LI6251); both are anti-correlated with
			  temperature


112: Status, Site , Mon 14-Jan-2002 11:47:17 MST, Daily Status, 1/14/02
FLOSS daily status, 1/14/02

Looked at past 72 hours of data

Meteorology
-----------
Prop Spd and Dir:	ok, winds mostly 180-270 degrees
T/RH (ISFF):		ok; did not examine thermocouple data
Pressure:		tower 0.5-2.5 mb higher than bare/sage;
			  sage 0.1 mb higher than bare
Rainr:			43 mm precip past 72 hours
Depth:			"snow depth" ~ 2050 mm; 
			  not much correlation apparent with precip
Rsw:			tower & sage only; albedo decreasing at tower
Rlw:			tower & sage only; ok
Rnet:			ok

Soil
----
Tsoil:			sage has a little larger diurnal cycle than
			  tower or bare
Gsoil:			sage has smallest diurnal cycle, bare greatest
Qsoil:			sage driest, bare wettest

Sonics
------
spd & dir:		ok
variances:		ok
u*, sigma_w/u*, w'tc':	ok

Scalars
-------
w'h2o':			KH2O ok; correlation coefficient between 
 			  LI7500 and KH2O is 0.97; linear fit is 
			  w'h2o'(LI7500) = w'h2o'(KH2O)*36.9 + 0.0096
w'co2':			not much correlation evident between w'co2' 
			  from LI7500 and w'co2mr' from LI6251
co2:			correlation coefficient between co2(LI7500) and 
                          co2mr (LI6251) is 0.93; 
			  linear fit is co2mr = 33.3*co2mr - 16.3



79: Status, Site all stns, Sun 09-Dec-2001 13:28:38 MST, visitors
Bob Kelly (Univ. of Wyoming) was here for a tour this morning.

Other visitors we've had:

Ed - (the local NWS cooperative observer for the past 25 years) was here for
a tour Friday afternoon.

Todd S- and Pam - from the Wildlife Refuge were here Thursday to teach Larry
and Dean about local plants.

Mark Lanier from the Wildlife Refuge was here Wed? last week to see what we
were up to.

We've (mostly Larry) also fielded reporter phone calls from the Denver Post 
and Colorado Springs Gazette last week.

We're expecting a visit by Matt Landis and other local High School teachers
tomorrow afternoon.

76: Status, Site all stns, Sat 08-Dec-2001 17:29:35 MST, shallow inversion today
We noticed a very shallow inversion this afternoon, where the max temperature
was at 5m.  The temperature differences were about 0.2 C from 0.5m to 5m and
back from 5m to 20m.  I believe the temperature sensors by now.  Later this
evening, it broke down and became completely stable throughout the profile.

70: Status, Site all stns, Sat 08-Dec-2001 10:12:55 MST, serial numbers and heights
These are the serial numbers for sensors that I know of:

Main:	TRH	Prop	CSAT	KH20	Baro	Logger
20	704	0007	0540	1394
15	701	0006
10	201	0003	0537	1393
5	001	0002	
2	004	0001	0536	
1	702	0004
0.5	005
Bare:
2	007		0539	1392	1002	1

Note that I'm just about to change the TRH at 20m to SN 104. (Done 10:55.)

Here are the heights of these sensors:

TRH
20	20.40+0.31-0.04-0.57 = 20.10
15	16.72-0.72-0.04-0.57 = 15.39
10	11.20-0.67-0.04-0.57 = 9.92
5	5.78+0.31-0.04-0.57 = 5.48
2	2.10+0.46-0.04-0.57 = 1.95
1	2.10-0.55-0.04-0.57 = 0.94
0.5	2.10-1.02-0.04-0.57 = 0.47

Prop
20	20.40-0.84-0.06+0.45 = 19.95
15	14.88-0.01+0.45 = 15.32
10	9.46+0.05-0.06+0.45 = 9.90
5	5.78-0.86-0.06+0.45 = 5.31
2	2.10-0.51-0.06+0.45 = 1.98
1	2.10-1.57-0.06+0.45 = 0.92

CSAT
20	20.40-0.89-0.06+0.05 = 19.50
10	11.20-0.89-0.06+0.05 = 10.30
2	2.10+0.04+0.05 = 2.19

KH20 relative to CSAT: 20cm down, 35cm back
LI6251 inlet frit relative to CSAT: 10cm down, 33cm back
LI7500 center of path relative to CSAT: 40cm down, 67cm back





		20.40-0.84-0.06+0.45
51: Status, Site tower, Wed 05-Dec-2001 12:54:23 MST, site visit
I spent the last 20 minutes at the main tower for daily stuff:
- dumped water in rain gauge (appeared to work!)
- shot 3 photos
- measured TRIME moisture 6.9 -> 7.4%

Now off to bare site for TRIME stuff and change wind gen pot.

47: Status, Site base, Tue 04-Dec-2001 21:34:38 MST, second mouse trapped
After seeing it scurrying across the back of the trailer today, I moved the
trap tonight.  It was filled 20 minutes later.  10 seconds of silence, everyone.

37: Status, Site base, Mon 03-Dec-2001 18:04:08 MST, covars restarted
I just added X&Ylevels to covars and restarted.

31: Status, Site stn2, Mon 03-Dec-2001 10:12:42 MST, misc
I just measured the sonic height as:
79" + 2.5" + 1.2" = 82.7" = 2.10m

I also took several photos from NW-W-SE at about 10:00

I also just measured power.  Solar, battery about 14.16V.  Wind generator 
by itself (disconnected) about 16.86V.  When the wind generator is plugged in,
it runs significantly slower.  This doesn't seem right.  When the battery is
low (and solar down), would the generator run even slower?

29: Status, Site all stns, Mon 03-Dec-2001 08:46:28 MST, daily status
Daily status:

Weather: almost completely clear (some clouds over mountains), snow disappearing
(10-20% bare?).  I'll try to take tower photos today.

cosmos-up
marigold-up (goldenrod hasn't crashed since replaced)
stn2-up
aster-gives RPC error

main site:
props    all okay
T/RH     something has been weird at 15 and/or 20m.  Maybe I'll just swap these.
Tc       5.75m died.  Others okay?  I'll have to make a daily plot that adds in
         the reference.  The 21x values appear to be somewhat noisy.
sonics   okay, though more spikes than I'd like (20-30 overnight at all levels)
kryptons okay, the 2m sensors have had spikes (large variance) in the data.
         I'll clean these soon, though the spikes are probably from blowing snow
licorCO2 okay, cal gases are cycling
rad	 okay I guess.  Haven't looked at daily plots yet.
soil	 okay I guess.  "
icedet   no action (which is good)
levels   okay
baro     okay
precip   okay?  reading zero.  I should do a calibration.
snowdepth okay?  may be seeing blowing grasses
power	 no problems

bare site:
T/RH	 okay?  The fan acted strange last night, perhaps due to power
sonic	 good.  The fewest spikes (perhaps the least blowing snow?)
krypton  okay.  weird values on cockpit probably from my work last night
heater   false high readings.  need to investigate this.
rad      okay.  need to look at daily plots
soil     okay.  "
levels   weird excursions.  need to check this
baro     okay
gps      okay
power    mostly only charging during day -- problem with wind generator?

need to rabbit-proof all cables!




25: Status, Site all stns, Sun 02-Dec-2001 08:33:21 MST, Dec 2 morning status
Weather:
  light snow last night, windy.  Small 8" drifts on refuge road.


Data system:
  up.  Marigold/arnica made it through the night!

Sonics:
  
  some non-zero diag values on all sonics. A few values >=16 which
    indicate a lost sample.
  all sonics also show a few scattered non-zero flags

  heater value at bare site always on

Kryptons: OK
  
TRH: OK

Props: OK

Thermocouples:  Sean had 4 deployed last night.  2 of the 4 are no
  longer reporting, assumed broken.

Radiation: values appear to be in acceptable range - did not do
  a thorough check.  
  Xlevel.bare and Ylevel.bare show some values around 10.

PAM power system:  still some question about whether wind generator
  voltage is getting to charging system.
  Current values: 
	Vbatt: 12.4 V
 	Icharge: 1.5 - 2.6 A  (sun through thin clouds, 7 m/s winds)
	Iload: 1.7 A
	

22: Status, Site all stns, Sat 01-Dec-2001 16:23:38 MST, Operational!

All NCAR sensors deployed and reporting data, however quality
checks have been minimal.

Prop angles have been entered, so prop wind directions should be
correct.

Licor co2 is deployed.  See other entry about calibration gas sequence.

Sean Burns has installed 4 (out of 20) thermocouples on the tower.
They are at 14, 15.5, 17, and 18.5 meters.

Kurt, SteveS and Tony left at about 3:00 this afternoon.  SteveO due
back tomorrow.


14: Status, Site all stns, Thu 29-Nov-2001 17:49:08 MST, status
Today's work:
- installed all T/RHs on main
- intalled all props on main
- installed 2m CSAT & KH20 on main
- set up rad stand (but not installed in final position)
- set up bare site infrastructure and some sensors

Other:
- Larry Mahrt and Dean Vickers arrived, staked out their sites and organized their equipment
- Sean Burns arrived, will set up tomorrow
- Mark Lanier stopped by for a visit

10: Status, Site tower, Thu 29-Nov-2001 02:25:48 GMT, Tower heights
I crudely set heights of booms today.  Here is what I did (NOTE THE FUTURE WORK LISTED IN 7-9):
1. The wood beams for the tower base are 10cm high
2. The metal plates under the jacks are ~1.5cm thick (probably 1/2")
3. The height of the top of the tube at the top of the first ladder above the metal plates for each of the 4 legs were: 199, 196, 197, & 202 cm
4. I average the values in 3. and add to the values of 1. and 2. to get a nominal height above ground of the top of the first ladder of 210 cm
5. We measured the height of one section as 184cm
6. Thus, I get the following heights above ground of the top of the tube at the top of each of the ladder sections (in cm):
210
394
578
762
946
1120
1304
1488
1672
1856
2040
2224
2408
These are labeled using duct tape on the tubes.
7. We then attempted to position each of the sensors as close as possible to the specified sensor heights, 0.5, 1, 2, 5, 10, 15, & 20m.  For the sensors up to 2m, we got to within 10cm by special clamping (still not perfect due to conflicts with brackets, etc.  For the sensors above 2m, we should be within 50cm.  We assumed that the sonics were 5cm above the boom, the props 50cm above the boom, and the T/RHs 50cm below the boom.  Also, anemomber yokes make the boom height ~6cm lower than the tube it is clamped to and T/RH yokes (vertical) set the boom ~7cm from the bottom (or top) edge.  Note that CSATs end up with the array almost perfectly aligned with the center of the tube that the yoke is clamped to.  All these relative heights should be checked as well.
8. I'll have to measure the exact positions (still relative to the duct-tape labels) later.
9. Overall, I am reasonably comfortable with this approach.  However, the 2m sonic will end up closer to 2.5m due to the tube available to clamp to.  Also, the terrain to the south and west slope up a bit, so the sensors at the end of 3 and 4m long booms may be 10-15cm closer to the local surface.  I'll try to determine this as well at a later time.


9: Status, Site tower, Wed 28-Nov-2001 23:56:08 GMT, status
Stuff done Monday:
trailer didn't get here until 4:30.  powered up and porch set-up by 5:45.
(overcast, flurries, cold)

Stuff done Tuesday:
- power to tower
- phone connected
- base computers installed
- adams up and transmitting Tbox by ~5:00
(overcast, flurries, cold)

Stuff done today:
- hung all booms on tower
- connected T/RH at 0.5 and 1.0m
- rad stand erected near tower base, rads attached
John and Tony show up ~2:00
(clear, very cold -- -30 last night -- wind started up about 12:30)